OpenStudy (i_love_my_nieces):

Which ordered pair is NOT a solution of y = -12x - 3? 1. (-3, 33) 2. (8, -98) 3. (-9, 105) 4. (3, -39)

OpenStudy (anonymous):

2

OpenStudy (yanasidlinskiy):

Plug each of those in and see if they don't work

OpenStudy (nirmalnema):

put the givent values of x and if you dont get give value of y then simply its the ans

OpenStudy (anonymous):

the y value would be -99... not -98

OpenStudy (anonymous):

do you want me to show you how to do it?

OpenStudy (nirmalnema):

yes @clarkydances is correct

OpenStudy (yanasidlinskiy):

For example....You need the one that doesn't work. I'll show you how to do number one. 33=-12(-3) -3 33=36-3 33=33 ----Correct. So you need the one that is wrong. It can't be a because it is right. Do number two in the steps i showed you and someone will be glad to help you through:)

OpenStudy (anonymous):

... iwas going to explain it...but ok then

OpenStudy (yanasidlinskiy):

Sorry......

OpenStudy (i_love_my_nieces):

Thank's! 2?

OpenStudy (anonymous):

its ok lol...and yes 2